# AVHBench: A Cross-Modal Hallucination Benchmark for Audio-Visual Large Language Models

### v1

@@ Appendix A Data samples and source code
Appendix B Additional results
@@
Appendix C Details on the data construction pipeline and statistics
@@ C.2 Positive and negative samples in QnAs
C.3 Details on human annotation
@@
C.4 Annotation quality control
@@
C.5 Sample outputs of each stage
Figure. S6 shows output samples from both Stage 1 and Stage 2 after automatic pipeline and human verification. Given the input audio-visual signals, such as audio (or audio-visual) captions and visual taggings, Stage 1 outputs the disentangeld audio and visual information. Using this disentangled information, Stage 2 outputs QnA pairs for judgment tasks and audio-visual captions for the description task.
C.6 Dataset statistic
In the main text, our proposed AVHBench dataset encompasses a total of 2,327 videos, comprising 1,238 real and 1,089 synthetic videos. The total QnA pairs are 5,816, distributed as follows: 1,250 for Audio-driven Video Hallucination, 2,508 for Video-driven Audio Hallucination, and 2,058 for Audio-visual Matching. In addition, the total audio-visual captions are 1,238. We have kept expanding our dataset during the submission. Our up-to-date dataset now includes 7,327 in-the-wild videos, along with 18,314 QnA pairs and 6,243 audio-visual captions. The QnA pairs include 3,936 pairs for the Audio-driven Video Hallucination task, 7,898 pairs for the Video-driven Audio Hallucination task, and 6,480 pairs for the Audio-visual Matching task. Therefore, the experiments conducted and the statistics summarized in the main text represent a subset of our full benchmark dataset.
@@ C.6 Dataset statistic
Appendix D Details on the evaluation settings
@@
D.1 Answer parsing strategy
@@ D.2 Benchmarked models
Appendix E Training details on Video-LLaMA audio alignment and LoRA fine-tuning
@@
To enhance the alignment between audio features and LLM, we pre-train Q-Formert (Li et al., 2023a) and a linear layer that link the frozen audio encoder and LLM, while keeping the visual-text branch and the LLM fixed. We train the audio-text branch of Video-LLaMA using total 173K audio-text pairs collected from AudioCaps (Kim et al., 2019) (50K), Clotho (Drossos et al., 2020) (14K), Audioset-SL (Hershey et al., 2021) (108K) and SoundBible (Mei et al., 2023) (1K). The training objective is to minimize the next token prediction loss (i.e., cross-entropy loss) over the textual responses, enhancing the model’s capability to understand and map audio data to textual representations. We utilize 4 A6000 (48GB) for distributed traning with batch size 32 per device and an initial learning rate (3e-5) and weight decay (0.05) for 1 epoch. We leverage mixed precision that uses fp16 for multiplication and fp32 for addition.
@@
We fine-tune Video-LLaMA using the training split of our AVHBench dataset, which consists of 10,327 videos with 87,624 QnA pairs. The training dataset is also constructed through the automatic pipeline outlined in Secion. 3.2 of the main paper, without any human verification process. We apply LoRA to the query, key, value, and output projections of Vicuna-7B, while the other parameters are frozen. We set the rank and alpha value of LoRA to 16 and 32, respectively. The training objective remains the same as the audio alignment stage. We utilize 4 NVIDIA A6000 (48GB) for distributed training with batch size 8 per device and an initial learning rate (3e-5) and weight decay (0.05) for 1 epoch. We also leverage mixed precision that uses fp16 for multiplication and fp32 for addition.
Appendix F GPT4 prompts for GAVIE metric

### v2

@@ Appendix A Data samples and source code
Appendix B Additional qualitative results
@@
Table S1: Additional evaluation results on AVHBench. We evaluate various more recent audio-visual LLMs and uinmodal LLMs on our proposed AVHBench. In the Audio-driven Video and Video-driven Audio Hallucination, both audio-visual signals are used as inputs for the model. For Video and Audio Hallucinations, only video or audio inputs are used, respectively. Acc. denotes the accuracy, Yes (%) is the proportion of “Yes” answers among total responses, and bold numbers stand for the best.
Video-LLaMA 50.1 50.1 100 66.7 99.9 50.2 50.2 100 66.9 100
Video-SALMONN 78.1 74.9 84.5 79.4 56.4 65.2 62.3 76.9 68.8 61.7
Video-LLaMA2 75.2 73.6 78.7 76.1 53.6 74.2 69.4 86.6 77.0 62.4
Gemini-Flash 83.3 85.7 81.0 83.7 47.3 63.0 57.9 94.7 71.9 81.7
Model Video Hallucination Audio Hallucination
Video-LLaMA 50.0 50.0 100 66.7 100 58.4 55.7 93.3 69.8 86.2
Video-SALMONN 82.3 83.9 79.9 81.9 47.6 66.5 70.5 56.8 62.9 40.3
Video-LLaMA2 80.7 87.0 72.1 78.8 41.4 79.8 81.9 76.4 79.1 46.7
Gemini-Flash 84.3 81.6 72.2 76.6 31.4 65.2 59.6 94.3 73.0 79.1
PLLaVA 81.5 77.9 87.8 82.6 56.3 - - - - -
LLaVA-OneVision 84.0 90.9 75.5 82.5 41.5 - - - - -
LTU - - - - - 67.2 63.9 84.2 72.7 68.3
Qwen2-Audio - - - - - 74.4 81.3 68.8 74.5 13.6
Appendix C Additional quantitative results
We conduct further evaluations on AVHBench using more recent audio-visual LLMs, such as Video-SALMONN (Sun et al., 2024), Video-LLaMA2 (Cheng et al., 2024), and GEMINI-Flash (Anil et al., 2023), and unimodal LLMs, such as LLaVA-OneVision (Li et al., 2024) and Qwen2-Audio (Chu et al., 2024). The results are summarized in Table S1, and we also include reference results from Video-LLaMA (Zhang et al., 2023a), PLLaVA (Xu et al., 2024), and LTU (Gong et al., 2024). We generally observe that recent models are improving on our benchmark, for both audio-visual LLMs and unimodal LLMs. GEMINI significantly outperforms other audio-visual LLMs in Audio-driven Video Hallucination. In Video-driven Audio Hallucination, Video-LLaMA2 exhibits strong performance compared to other models. Interestingly, we find that overall model performance improves when using unimodal signals for evaluation as summarized in Video and Audio Hallucination results. Specifically, Video-LLaMA2 demonstrates notable improvements with unimodal signals compared to using both audio-visual signals, supporting discussions in Sec. 4.2 on how cross-modal signals can confuse models and evoke hallucinations. These findings highlight recent enhancements in models to address cross-modal hallucination and also describe the effectiveness of our benchmark in evaluating these advancements.
Appendix D Details on the data construction pipeline and statistics
@@ D.1 GPT4 prompts
D.2 Positive and negative samples in QnAs
D.3 Details on human annotation
@@
D.4 Annotation quality control
@@
(a) A sample of error correction for Stage 1 outputs
(b) A sample of error correction for Stage 2 outputs
Figure S6: Samples of model-based annotation error correction in the data construction pipeline. (a) Although the automatic disentanglement in Stage 1 may include noisy annotations due to reliance on off-the-shelf models, we observe that human annotators effectively identify and correct these inaccuracies. Similarly, (b) The initial audio-visual captions in Stage 2 may inaccurately describe the given video, as ChatGPT utilizes unaligned audio-only and video-only captions that may contain errors for generating audio-visual captions. However, human verification plays a crucial role in recognizing and correcting these errors.
@@
D.5 Model-based annotation error study
As automatic annotations rely on off-the-shelf models, initial errors are unavoidable. To address this, human verification plays a crucial role in our semi-automatic dataset construction pipeline for correcting errors that sometimes occur from noisy model outputs. We provide examples of how this process enhances annotation accuracy. For example, in Stage 1, RAM+++++ + (Zhang et al., 2023c), which extracts visual tags from videos, might include noisy tags that misrepresent the audio-visual information. As shown in Fig. S6-(a), despite the absence of a “plane” in the video, it is listed as an in-view silent object because it was mistakenly tagged by RAM+++++ +. However, we observe that Stage 1 annotators successfully identify and remove these inaccuracies during the annotation process.
Moreover, similar errors can occur in Stage 2, when ChatGPT uses audio-only and video-only captions to generate an audio-visual caption. Since captions from each modality could be misaligned, the language model (ChatGPT) may not always accurately capture their correspondence, leading to merging these misaligned audio-visual captions. For example, as shown in Fig. S6-(b), although the video caption incorrectly describes a man sitting on the rock instead of a woman, the initial audio-visual caption generated by ChatGPT merely merges the two captions from different modalities, regardless of the misalignment, resulting in an incorrect audio-visual caption. Nevertheless, during human verification, the error is identified, and the caption is modified from “with a man” to “and she is,” ensuring the refined audio-visual caption accurately describes the video.
D.6 Sample outputs of each stage
Figure. S7 shows output samples from both Stage 1 and Stage 2 after automatic pipeline and human verification. Given the input audio-visual signals, such as audio (or audio-visual) captions and visual taggings, Stage 1 outputs the disentangled audio and visual information. Using this disentangled information, Stage 2 outputs QnA pairs for judgment tasks and audio-visual captions for the description task.
D.7 Dataset statistic
@@
Appendix E Details on the evaluation settings
@@
E.1 Answer parsing strategy
@@ E.2 Benchmarked models
Appendix F Training details on Video-LLaMA audio alignment and LoRA fine-tuning
@@
To enhance the alignment between audio features and LLM, we pre-train Q-Former (Li et al., 2023a) and a linear layer that link the frozen audio encoder and LLM, while keeping the visual-text branch and the LLM fixed. We train the audio-text branch of Video-LLaMA using total 173K audio-text pairs collected from AudioCaps (Kim et al., 2019) (50K), Clotho (Drossos et al., 2020) (14K), Audioset-SL (Hershey et al., 2021) (108K) and SoundBible (Mei et al., 2023) (1K). The training objective is to minimize the next token prediction loss (i.e., cross-entropy loss) over the textual responses, enhancing the model’s capability to understand and map audio data to textual representations. We utilize 4 A6000 (48GB) for distributed training with batch size 32 per device and an initial learning rate (3e-5) and weight decay (0.05) for 1 epoch. We leverage mixed precision that uses fp16 for multiplication and fp32 for addition.
@@
We fine-tune Video-LLaMA using the training split of our AVHBench dataset, which consists of 10,327 videos with 87,624 QnA pairs. The training dataset is also constructed through the automatic pipeline outlined in Section. 3.2 of the main paper, without any human verification process. We apply LoRA to the query, key, value, and output projections of Vicuna-7B, while the other parameters are frozen. We set the rank and alpha value of LoRA to 16 and 32, respectively. The training objective remains the same as the audio alignment stage. We utilize 4 NVIDIA A6000 (48GB) for distributed training with batch size 8 per device and an initial learning rate (3e-5) and weight decay (0.05) for 1 epoch. We also leverage mixed precision that uses fp16 for multiplication and fp32 for addition.
Appendix G GPT4 prompts for GAVIE metric
@@
Appendix H Maintenance
Figure S9: Dataset maintenance. We plan to maintain a benchmark leaderboard webpage to facilitate ongoing evaluation of audio-visual LLMs.
To further support ongoing research on audio-visual LLMs, we plan to maintain a benchmark leaderboard webpage as shown in Fig. S9. The leaderboard will list models we have already evaluated and will be continuously updated as new audio-visual LLMs are tested on our benchmark.